OpenStudy (anonymous):

LAST QUESTION.....The position of an object moving in a straight line is given by s=2t^2-3t, where s is the distance in meters and t is the time in seconds the object has been in motion, How long,(to the nearest tenth) will it take the object to move 14 meters?

OpenStudy (anonymous):

is t=7/2

OpenStudy (anonymous):

So\[14=2t^2-3t \rightarrow 2t^2-3t-14=0\rightarrow 2t^2-7t+4t-14=0\rightarrow \]\[(t+2)(2t-7)=0\rightarrow t=-2 or t=7/2\]Since the model does not apply for times less than zero, the solution is t=7/2.

OpenStudy (anonymous):

Sorry one more then im done, A merchant has coffee worth $20 a pound that she wished to mix with 70 pounds of coffee workth $80 a pound to get a mixture that can be sold for $40 a pound. How many pounds of the $20 coffee should be used??

OpenStudy (anonymous):

20x + (80)(70) = 40 (70+x)

OpenStudy (anonymous):

2800 = 20x 140 = x

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No sweat.